# Strict quantization for compact pseudo-Kähler manifolds and group actions

### v1

@@ 3.1 Vector valued reproducing kernel
Here, we recall some definitions concerning reproducing kernel Hilbert spaces for vector valued functions, we refer to [BM] and [Pe] (see also [CVT, Section 2]).
@@
Let $X$ be a manifold and $E$ be a vector bundle on $X$. A $E$-valued reproducing kernel Hilbert space on $X$ is a Hilbert space $H$ such that:
@@
    for all $x\in X$ there exists a positive constant $C_{x}$ such that
    $\lVert s(x)\rVert_{E_{x}}\leq C_{x}\lVert s\rVert_{H}\,,$
    for each $s\in H$.
@@
$\mathrm{ev}_{x}\,:\,H^{G}_{k}(X)\rightarrow T^{*\,(0,q)}X,\qquad\mathrm{ev}_{x}(s)=s(x)$
is a bounded operator and the reproducing kernel associated with $H^{G}_{k}(X)$ is $S^{G}$. The kernel $S^{G}$ reproduces the value of a section $s\in H^{G}_{k}(X)$ at point $x\in X$. Indeed, for all $x\in X$ and $v\in T^{*\,(0,q)}_{x}X$
$\mathrm{ev}_{x}^{*}\,v=S^{G}_{k}(\cdot,\,x)\,v$
so that
$\langle s(x)|v\rangle=(s,\,S^{G}_{k}(\cdot,\,x)v)\,.$
Thus, for each $x\in X$ and $v\in T^{*\,(0,q)}_{x}X$ the $(0,q)$-forms $s_{x,v}=S^{G}_{k}(\cdot,\,x)\,v$ is called the coherent state (of level $k$) associated to $x$ and $v$.
@@
$\sigma_{v}^{(k)}(A)\,:\,X\rightarrow\mathbb{C},\qquad x\mapsto\sigma_{v}(A)(x):=\frac{\left(s_{x,v}|\,As_{x,v}\right)}{\left(s_{x,v}|\,s_{x,v}\right)}$
@@
The definition of coherent states goes back to Berezin. A coordinate independent version and extensions to line bundles were given by Rawnsley [R], we mainly refer to [S2] where the coherent vectors are parameterized by the elements of $L^{*}\setminus 0$, where $L$ is a positive line bundle of a given Hodge manifold. In the following we shall generalize the coherent states in our setting.
Let $f$ be a $G$ invariant function on $\mu^{-1}(0)$, for any extension $\widetilde{f}$ on $X$, we can assign to it its Toeplitz operator $T^{G}_{k}[\widetilde{f}]$ and then assign to it the covariant symbol, as in Definition 3.2
$\sigma_{v}^{(k)}\left(T^{G}_{k}[\widetilde{f}]\right)\,:\,X\rightarrow\mathbb{C}$
@@ 3.2 Berezin transforms for Toeplitz operators
Recall that $\pi\,:\,\mu^{-1}(0)\rightarrow X_{G}$ is the projection. Given a smooth function $f\in C^{\infty}(X_{G})$ invariant under the circle action, we define:
@@
The map
$C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}\,,\qquad f\mapsto I_{k}^{G}[f]:=\sigma^{(k)}_{v}(T^{G}_{k}[f])$
is called the Berezin transform of level $k$.
@@
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, there exist local coordinates $(x_{1},\dots\,x_{2n+1})$ on $U$ such that the Berezin transform $I^{G}_{k}[f]$ evaluated at the point $x\in U$ has a complete asymptotic expansion in decreasing integer powers of $k$
@@
as $k$ goes to infinity, where $I_{j}\,:\,C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}$ are maps such that $I_{0}[f]=f$.
Proof of Theorem 3.1.
Let $f\in C^{\infty}(X_{G})^{S^{1}}$ and let us first study the numerator of
$\sigma^{(k)}_{v}(T^{G}_{k}[\widetilde{f}])$
which is
$\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[\widetilde{f}]\,S^{(q)}_{k}(\cdot,\,x)\,v\right)=\int_{X}v^{\dagger}\,S^{(q)}_{k}(x,\,y)\widetilde{f}(y)\,S^{(q)}_{k}(y,\,x)\,v\,\mathrm{dV}_{X}(y)\,.$
@@
$\left(S^{(q)}_{k}(\cdot,\,x)\,v\bigg{|}T^{(q)}_{k}[\widetilde{f}]\,S^{(q)}_{k}(\cdot,\,x)\,v\right)\sim k^{d}\,v^{\dagger}\sigma^{0}_{T^{G}[\widetilde{f}]}(x,\,x)\,v+O(k^{d-1})$
@@
$\sigma^{0}_{T^{G}[\widetilde{f}]}(x,x)=\frac{1}{2}\pi^{-n-1}\left|{\rm det\,}L_{x}\right|\,\widetilde{f}(x)\,\tau_{x,n_{-}}\,,$
for every $x\in\mu^{-1}(0)$. Thus, taking the quotient we get
$I_{k}^{G}[\widetilde{f}](x)=\widetilde{f}(x)+O(k^{d-1})\,$
for each $x\in\mu^{-1}(0)$. ∎
Theorem 3.1 is consequence of the asymptotic expansion of Toeplitz operators for $(0,q)$-forms and the stationary phase formula. Here, we shall mainly refer to [GH1] where we study Toeplitz operators on CR manifolds, see also [MM2] and [Hs] for the asymptotic behavior of the Bergman kernel for $(0,q)$-forms, and references therein.
4 Quotients and complexification of groups
Here, we recall some results proved in [GS, Section 4] and we explain how to extend it to pseudo Kähler manifolds. There exists a unique connected complex Lie group, $G^{\mathbb{C}}$, with the following two properties:
@@ 4 Quotients and complexification of groups
Now, we shall prove some properties of $G$-actions on pseudo-Kähler manifolds. Now let $M$ be a $2n$-dimensional symplectic manifold. A polarization of $X$ is an integrable Lagrangian subbundle, $F$, of $TM\otimes\mathbb{C}$. A symplectic manifold is a pseudo-Kähler manifold if it possesses a polarization $F$. The next lemma, an analogue of [GS, Lemma 4.3], will be used to reconcile this definition with the standard one:
@@
Let $V$ be a (real) symplectic vector space with symplectic form, $\Omega$. Let $F$ be a Lagrangian subspace of $V\otimes\mathbb{C}$ Then there exists a unique linear mapping $J\,:\,V\rightarrow V$ such that
  • i)
    $J^{2}=-I$,
  • ii)
    $F=\{v+\sqrt{-1}\,Jv,\,v\in V\}$,
  • iii)
    $\Omega(Jv,\,Jw)=\Omega(v,\,w)$,
  • iv)
    The quadratic form
    $B(v,\,w)=\Omega(v,Jw)$
    is symmetric and non-degenerate.
Let $M$ be a symplectic manifold and $F$ a polarization. By the lemma we get for each $m\in M$ a mapping $J_{m}\,:\,T_{m}M\rightarrow T_{m}M$ with the properties i), ii), and iii) and a quadratic form, $B_{m}$, on $M$. Furthermore $J$ defines an almost-complex structure on $M$ and $B$ a pseudo-Riemannian structure. The integrability of $F$ implies that the almost-complex structure is complex. Therefore, the quadruple $(M,J,B,\Omega)$ is pseudo-Kähler manifold in the usual sense.
Let $(M,F)$ be a compact Kähler manifold and $G$ a compact connected Lie group which acts on $M$, preserving $F$. We shall prove Theorem 4.1 below, which is an analogue of [GS, Theorem 4.4] and in fact the first part of the proof is identical. However, note that [GS, Theorem 4.4] relies on the fact that the group of analytic diffeomorphisms of $M$ which preserve $F$ is a finite dimensional Lie group, see [K]. Here, we prove that the former result still holds true in the setting of pseudo-Riemannian manifolds (in Theorem 4.1 we assume that $G^{\mathbb{C}}$ is simply connected, for the general case we refer to [GS, Theorem 4.4]).
Let $\mathrm{Iso}(X,\,g)$ be the group of isometries from $(M,\,g)$ onto itself, that is $f\in\mathrm{Iso}(X,\,g)$ if and only if $f$ is a smooth diffeomorphism and $f^{*}g=g$. Let $(M,\,T^{1,0}M)$ be a connected complex manifold and denote by $\mathrm{Iso}(M,\,g)$ the group of isometries on $M$ with respect to some Riemannian metric $g$. Let $\mathrm{Aut}_{J}(M)$ be the group of complex automorphisms on $M$, that is $f\in\mathrm{Aut}_{J}(M)$ if and only if $f\,:\,M\rightarrow M$ is a smooth diffeomorphism satisfying $\mathrm{d}f(T^{1,0}M)\subseteq T^{1,0}M$.
@@
Let us assume that $G^{\mathbb{C}}$ is simply-connected. The action of $G$ can be canonically extended to an action of $G^{\mathbb{C}}$, preserving the pseudo- Kähler structure of $M$.
@@
Let
$\tau\,:\,\mathfrak{g}^{\mathbb{C}}\rightarrow\text{(real) vector fields on $M$}$
be the mapping,
$\xi^{(1)}+\sqrt{-1}\,\xi^{(2)}\mapsto\xi^{(1)}_{M}+J\,\xi^{(2)}_{M}$
where $\xi^{(1)},\,\xi^{(2)}\in\mathfrak{g}$. By (4.2) and (4.3) in [GS], $\tau$ is a morphism of Lie algebras. Moreover, by (4.2), if $\eta\in\mathfrak{g}^{\mathbb{C}}$, $\tau(\eta)$ is a vector field preserving $F$. By Lemma 4.2 below, $\mathrm{Iso}(M,\,g)\cap\mathrm{Aut}_{J}(M)$ is a (finite dimensional) Lie group; therefore, if $G^{\mathbb{C}}$ is simply-connected, $\tau$ can be extended uniquely to a morphism of Lie groups. ∎
@@
By [Ph] and [K], we have that $\mathrm{Iso}(M,\,g)$ is a finite-dimensional Lie group, since $(M,\,g)$ is a pseudo-Riemannian manifold. Recall that a subgroup of a Lie group is itself a Lie group if it is a closed subgroup.
Now, consider the subgroup $\mathrm{Iso}(M,\,g,\,J)\subset\mathrm{Iso}(M,\,g)$, consisting of those diffeomorphisms that preserve both the metric $g$ and the complex structure $J$. The preservation of $g$ and $J$ by a diffeomorphism are smooth conditions. Thus, we can consider the set $\mathrm{Iso}(M,\,g,\,J)$ as the solution set of these equations within the group $\mathrm{Iso}(M,\,g)$. If a sequence of diffeomorphisms in $\mathrm{Iso}(M,\,g,\,J)$ converges to a diffeomorphism in the topology of $\mathrm{Iso}(M,\,g)$ (which is the compact-open topology), the limit will also preserve $g$ and $J$.
@@
Recall now that $(M,\,\omega)$ is the complex pseudo-Kähler manifold obtained by quotienting out by the circle action on $X$, $M:=X/S^{1}$. The CR action of $G$ on $X$ descends to a complex and Hamiltonian action on $M$ with moment map $\Phi\,:\,M\rightarrow\mathfrak{g}^{*}$. Thus, we can form the reduced space
$M_{G}:=\Phi^{-1}(0)/G\,.$
Let $M_{s}$ be the saturation of $\Phi^{-1}(0)$ with respect to $G^{\mathbb{C}}$,
$M_{s}:=\{g\circ x\,:\,x\in\Phi^{-1}(0),\,g\in G^{\mathbb{C}}\}\,,$
the points of $M_{s}$ are called stable points. In the same way as in [GS, Theorem 4.5], one can prove the following theorem. Here, we shall retrace the proof again since the assumption (4) will be crucial in the pseudo-Riemannian case.
@@
The set $M_{s}$ is an open subset of $M$ and $G^{\mathbb{C}}$ acts freely on it. Then, $M_{G}$ can be represented as the quotient space $M_{G}:=M_{s}/G^{\mathbb{C}}$.
@@
Recall that $b$ is the nondegenerate bilinear form on $HX$ such that
$b(\cdot,\cdot)=\mathrm{d}\omega_{0}(\cdot,J\cdot)\,.$
@@
Let us denote $\mu^{-1}(0)$ by $Y$. There is a natural projection $\pi_{Y}\,:\,Y\rightarrow\Phi^{-1}(0)$. Recall that we assume that 00 is a regular value of $\mu$, and that the action of $G$ on $\mu^{-1}(0)$ is free and
[…]

### v2

@@ 4.1 Vector valued reproducing kernel
Here, we recall some definitions concerning reproducing kernel Hilbert spaces for vector valued functions, we refer to [3, 28] (see also [4, Section 2]).
@@
Let $X$ be a manifold and $E$ be a vector bundle on $X$. A $E$-valued reproducing kernel Hilbert space on $X$ is a Hilbert space $H$ such that
@@
    for all $x\in X$ there exists a positive constant $C_{x}$ such that $\lVert s(x)\rVert_{E_{x}}\leq C_{x}\lVert s\rVert_{H}$ for each ${s\in H}$.
@@
$\mathrm{ev}_{x}\colon\ H^{G}_{k}(X)\rightarrow T^{*(0,q)}X,\qquad\mathrm{ev}_{x}(s)=s(x)$
is a bounded operator and the reproducing kernel associated with $H^{G}_{k}(X)$ is $S^{G}$. The kernel $S^{G}$ reproduces the value of a section $s\in H^{G}_{k}(X)$ at point $x\in X$. Indeed, for all $x\in X$ and $v\in T^{*(0,q)}_{x}X$ $\mathrm{ev}_{x}^{*}v=S^{G}_{k}(\cdot,x)v$ so that $\langle s(x)|v\rangle=\bigl{(}s,S^{G}_{k}(\cdot,x)v\bigr{)}$. Thus, for each $x\in X$ and $v\in T^{*(0,q)}_{x}X$ the $(0,q)$-forms $s_{x,v}=S^{G}_{k}(\cdot,x)v$ is called the coherent state (of level $k$) associated to $x$ and $v$.
@@
$\sigma_{v}^{(k)}(A)\colon\ X\rightarrow\mathbb{C},\qquad x\mapsto\sigma_{v}(A)(x):=\frac{(s_{x,v}|As_{x,v})}{(s_{x,v}|s_{x,v})}$
@@
The definition of coherent states goes back to Berezin. A coordinate independent version and extensions to line bundles were given by Rawnsley [30], we mainly refer to [33] where the coherent vectors are parameterized by the elements of $L^{*}\setminus 0$, where $L$ is a positive line bundle of a given Hodge manifold. In the following we shall generalize the coherent states in our setting.
Let $f$ be a $G$ invariant function on $\mu^{-1}(0)$, for any extension $\widetilde{f}$ on $X$, we assign to it its Toeplitz operator $T^{G}_{k}\big{[}\widetilde{f}\big{]}$ and then assign to it the covariant symbol, as in Definition 4.2
$\sigma_{v}^{(k)}\bigl{(}T^{G}_{k}\big{[}\widetilde{f}\big{]}\bigr{)}\colon\ X\rightarrow\mathbb{C}$
@@ 4.2 Berezin transforms for Toeplitz operators
Recall that $\pi\colon\mu^{-1}(0)\rightarrow X_{G}$ is the projection. Given a smooth function $f\in C^{\infty}(X_{G})$ invariant under the circle action, we define the Berezin transform of level $k$.
@@
The map $C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}$, $f\mapsto I_{k}^{G}[f]:=\sigma^{(k)}_{v}\bigl{(}T^{G}_{k}[f]\bigr{)}$ is called the Berezin transform of level $k$.
@@
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, there exist local coordinates $(x_{1},\dots,x_{2n+1})$ on $U$ such that the Berezin transform $I^{G}_{k}[f]$ evaluated at the point $x\in U$ has a complete asymptotic expansion in decreasing integer powers of $k$
@@
as $k$ goes to infinity, where $I_{j}\colon C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}$ are maps such that $I_{0}[f]=f$.
Proof.
Let $f\in C^{\infty}(X_{G})^{S^{1}}$ and let us first study the numerator of $\sigma^{(k)}_{v}\bigl{(}T^{G}_{k}\big{[}\widetilde{f}\big{]}\bigr{)}$ which is
$\bigl{(}S^{G}_{k}(\cdot,x)v\big{|}T^{G}_{k}\big{[}\widetilde{f}\big{]}S^{(q)}_{k}(\cdot,x)v\bigr{)}=\int_{X}v^{\dagger}S^{(q)}_{k}(x,y)\widetilde{f}(y)S^{(q)}_{k}(y,x)v\mathrm{dV}_{X}(y).$
@@
$\bigl{(}S^{(q)}_{k}(\cdot,x)v\big{|}T^{(q)}_{k}\big{[}\widetilde{f}\big{]}S^{(q)}_{k}(\cdot,x)v\bigr{)}\sim k^{d}v^{\dagger}\sigma^{0}_{T^{G}[\widetilde{f}]}(x,x)v+O\bigl{(}k^{d-1}\bigr{)}$
@@
$\sigma^{0}_{T^{G}[\widetilde{f}]}(x,x)=\frac{1}{2}\pi^{-n-1}\left|\det L_{x}\right|\widetilde{f}(x)\tau_{x,n_{-}},$
for every $x\in\mu^{-1}(0)$. Thus, taking the quotient we get $I_{k}^{G}\big{[}\widetilde{f}\big{]}(x)=\widetilde{f}(x)+O\bigl{(}k^{d-1}\bigr{)}$ for each ${x\in\mu^{-1}(0)}$. ∎
Theorem 4.5 is consequence of the asymptotic expansion of Toeplitz operators for $(0,q)$-forms and the stationary phase formula. Here, we shall mainly refer to [7] where we study Toeplitz operators on CR manifolds, see also [10, 23] for the asymptotic behavior of the Bergman kernel for $(0,q)$-forms, and references therein.
5 Quantization commutes with reduction
5.1 Quotients and complexification of groups
Let $G$ be a compact connected Lie group and let $\mathfrak{g}$ be its Lie algebra. Here, we recall some results proved in [9, Section 4] and we explain how to extend it to pseudo Kähler manifolds. There exists a unique connected complex Lie group, $G^{\mathbb{C}}$, with the following two properties:
@@ 5.1 Quotients and complexification of groups
Now, we shall prove some properties of $G$-actions on pseudo-Kähler manifolds. A polarization of $X$ is an integrable Lagrangian subbundle, $F$, of $TM\otimes\mathbb{C}$. The next lemma, an analogue of [9, Lemma 4.3], will be used to reconcile this definition with the standard one.
@@
Let $V$ be a ((((real)))) symplectic vector space with symplectic form, $\Omega$. Let $E$ be a Lagrangian subspace of $V\otimes\mathbb{C}$ Then there exists a unique linear mapping $K\colon V\rightarrow V$ such that
  • $(i)$
    $K^{2}=-I$,
  • $(ii)$
    $E=\big{\{}v+\sqrt{-1}Kv,v\in V\big{\}}$,
  • $(iii)$
    $\Omega(Kv,Kw)=\Omega(v,w)$,
  • $(iv)$
    The quadratic form $B(v,w)=\Omega(v,Kw)$ is symmetric and non-degenerate.
Let $(M,\omega)$ be a symplectic manifold and $F$ a polarization. By the lemma we get for each ${m\in M}$ a mapping $K=J_{m}\colon T_{m}M\rightarrow T_{m}M$ with the properties (i), (ii), and (iii) and a quadratic form, $B=g_{m}$, on $M$. Furthermore, $J$ defines an almost-complex structure on $M$ and $g$ a pseudo-Riemannian structure. The integrability of $F$ implies that the almost-complex structure is complex. Therefore, the quadruple $(M,J,g,\omega)$ is a pseudo-Kähler manifold in the usual sense.
Let $(M,\omega)$ be a compact pseudo-Kähler manifold and $G$ a compact connected Lie group which acts on $M$, preserving $F$. We shall prove Theorem 5.2 below, which is an analogue of [9, Theorem 4.4] and in fact the first part of the proof is identical. However, note that [9, Theorem 4.4] relies on the fact that the group of analytic diffeomorphisms of $M$ which preserve $F$ is a finite-dimensional Lie group, see [16]. Here, we prove that the former result still holds true in the setting of pseudo-Riemannian manifolds (in Theorem 5.2, we assume that $G^{\mathbb{C}}$ is simply connected, for the general case we refer to [9, Theorem 4.4]).
Let $\mathrm{Iso}(X,g)$ be the group of isometries from $(M,g)$ onto itself, that is $f\in\mathrm{Iso}(X,g)$ if and only if $f$ is a smooth diffeomorphism and $f^{*}g=g$. Let $\bigl{(}M,T^{1,0}M\bigr{)}$ be a connected complex manifold and denote by $\mathrm{Iso}(M,g)$ the group of isometries on $M$ with respect to some Riemannian metric $g$. Let $\mathrm{Aut}_{J}(M)$ be the group of complex automorphisms on $M$, that is $f\in\mathrm{Aut}_{J}(M)$ if and only if $f\colon M\rightarrow M$ is a smooth diffeomorphism satisfying ${\rm d}f\bigl{(}T^{1,0}M\bigr{)}\subseteq T^{1,0}M$.
@@
Let $G^{\mathbb{C}}$ be simply-connected. The action of $G$ can be canonically extended to an action of $G^{\mathbb{C}}$, preserving the pseudo-Kähler structure of $M$.
@@
Let $\tau\colon\mathfrak{g}^{\mathbb{C}}\rightarrow$ (real) vector fields on $M$ be the mapping, $\xi^{(1)}+\sqrt{-1}\xi^{(2)}\mapsto\xi^{(1)}_{M}+J\xi^{(2)}_{M}$ where $\xi^{(1)},\xi^{(2)}\in\mathfrak{g}$. By [9, equations (4.2) and (4.3)], $\tau$ is a morphism of Lie algebras. Moreover, by (4.2), if $\eta\in\mathfrak{g}^{\mathbb{C}}$, $\tau(\eta)$ is a vector field preserving $F$. By Lemma 5.3 below, $\mathrm{Iso}(M,g)\cap\mathrm{Aut}_{J}(M)$ is a (finite-dimensional) Lie group; therefore, if $G^{\mathbb{C}}$ is simply-connected, $\tau$ can be extended uniquely to a morphism of Lie groups. ∎
@@
By [16, 29], we have that $\mathrm{Iso}(M,g)$ is a finite-dimensional Lie group, since $(M,g)$ is a pseudo-Riemannian manifold. Recall that a subgroup of a Lie group is itself a Lie group if it is a closed subgroup.
Now, consider the subgroup $\mathrm{Iso}(M,g,J)\subset\mathrm{Iso}(M,g)$, consisting of those diffeomorphisms that preserve both the metric $g$ and the complex structure $J$. The preservation of $g$ and $J$ by a diffeomorphism are smooth conditions. Thus, we consider the set $\mathrm{Iso}(M,g,J)$ as the solution set of these equations within the group $\mathrm{Iso}(M,g)$. If a sequence of diffeomorphisms in $\mathrm{Iso}(M,g,J)$ converges to a diffeomorphism in the topology of $\mathrm{Iso}(M,g)$ (which is the compact-open topology), the limit will also preserve $g$ and $J$.
@@
Recall now that $(M,\omega)$ is the complex pseudo-Kähler manifold obtained by quotienting out by the circle action on $X$, $M:=X/S^{1}$. The CR action of $G$ on $X$ descends to a complex and Hamiltonian action on $M$ with moment map $\Phi\colon M\rightarrow\mathfrak{g}^{*}$. Thus, we form the reduced space ${M_{G}:=\Phi^{-1}(0)/G}$. Let $M_{s}$ be the saturation of $\Phi^{-1}(0)$ with respect to $G^{\mathbb{C}}$,
$M_{s}:=\big{\{}g\circ x\mid x\in\Phi^{-1}(0),\,g\in G^{\mathbb{C}}\big{\}},$
the points of $M_{s}$ are called stable points. In the same way as in [9, Theorem 4.5], one can prove the following theorem. Here, we shall retrace the proof again since we use the assumption that the pseudo-metric $g$ restricted to the orbit through any $m\in\Phi^{-1}(0)$ is non-degenerate.
@@
Assume that the pseudo-metric $g$ restricted to the orbit through any $m\in\Phi^{-1}(0)$ is non-degenerate and the action of $G$ on $\Phi^{-1}(0)$ is free. The set $M_{s}$ is an open subset of $M$ and $G^{\mathbb{C}}$ acts freely on it and $M_{G}$ can be represented as the quotient space $M_{G}:=M_{s}/G^{\mathbb{C}}$.
@@
Recall that $b$ is the nondegenerate bilinear form on $HX$ such that $b(\cdot,\cdot)={\rm d}\omega_{0}(\cdot,J\cdot)$. Let us denote $\mu^{-1}(0)$ by $Y$. There is a natural projection $\pi_{Y}\colon Y\rightarrow\Phi^{-1}(0)$. Recall that we assume that 00 is a regular value of $\mu$, and that the action of $G$ on $\mu^{-1}(0)$ is free. Furthermore, by hypothesis the bilinear form $b$ is nondegenerate on $Y$. Then, $\pi_{Y}(\underline{\mathfrak{g}}_{x})=T_{\pi(x)}\Phi^{-1}(0)$ and
$\big{\{}{\xi}_{M}(\pi(x))\in T_{\pi(x)}M\mid\xi\in\sqrt{-1}\mathfrak{g}\big{\}}$
$M_{s}=\bigcup_{g\in G^{\mathbb{C}}}gU,$
then $M_{s}$ is itself open. Thus, the stabilizer algebra of $x$ in $\mathfrak{g}^{\mathbb{C}}$ is zero; so the action of $G^{\mathbb{C}}$ on $M_{s}$ is locally free. To show that $G^{\mathbb{C}}$ acts freely on $M_{s}$ one can proceed as in [9, the second part of Theorem 4.5, p. 527]. ∎
5.2 Proof of Theorem 1.5
The action of $G$ on the line bundle, $L$, can be canonically extended to an action of $G^{\mathbb{C}}$ on $L$. The proof of this fact is identical with the proof of [9, Theorem 5.1], and we will omit it. Here, we just recall the infinitesimal action of an element $\eta=\sqrt{-1}\xi\in\sqrt{-1}\mathfrak{g}$ on a harmonic $L$-valued $(0,n_{-})$-form $s$
$-\nabla_{\eta_{M}}s-2\pi\langle\Phi(\cdot),{\xi}\rangle s.$ (5.1)
Let $s$ be a harmonic $L$-valued $(0,n_{-})$-form, and let $(s,s)(m)$ be the norm of $s$. By definition $(s,s)$ is a non-negative real-valued function. By assumption, (,)(\,,\,)( , ) is invariant with respect to parallel transport; so for all $\eta=\sqrt{-1}\xi\in\sqrt{-1}\mathfrak{g}$, $\eta_{M}(s,s)=(\nabla_{\eta_{M}}s,s)+(s,\nabla_{\eta_{M}}s)$. Suppose now that $s$ is $G^{\mathbb{C}}$-invariant. Then, by (5.1), $\nabla_{\eta_{M}}s=-2\pi\langle\Phi(\cdot),{\xi}\rangle s$ so $\eta_{M}\langle s,s\rangle=-4\pi\langle\Phi(\cdot),{\xi}\rangle\langle s,s\rangle$.
Now let $\mathcal{Q}(M)$ be the space of harmonic $L$-valued $(0,n_{-})$-forms and let $\mathcal{Q}(M_{s})$ be the space of harmonic $L$-valued $(0,n_{-})$-forms over $M_{s}$. Let $\mathcal{Q}(M)^{G}$ band $\mathcal{Q}(M_{s})^{G}$. be the set of G-fixed vectors in these two spaces. The following theorem follows as in [9, Theorem 5.2].
Theorem 5.5.
The canonical mapping $\mathcal{Q}(M_{s})^{G}\rightarrow\mathcal{Q}(M_{G})$ is bijective.
It is clear that the restriction mapping $\mathfrak{r}\colon\mathcal{Q}(M)^{G}\rightarrow\mathcal{Q}(M_{s})^{G}$ is injective; so, by Theorem 5.5, to prove “quantization commutes with reduction”, it is enough to prove that $\mathfrak{r}$ is surjective. Now, we shall need the following theorem.
Theorem 5.6.
The set $M\setminus M_{s}$ is contained in a complex sub-manifold of $M$ of complex codimension greater than one.
This theorem is an analogue of [9, Theorem 5.7] for pseudo-Riemannian manifolds. Note that here the existence theorem [9, Theorem 5.6] is not true in general for the pseudo-Riemannian case: we have to replace the spaces of smooth sections of $L^{k}$ with space of $L^{k}$-valued $(0,n_{-})$-forms, see Theorem 5.9. Then, Theorem 5.6 follows by combining Theorems 5.7 and 5.9 below.
Theorem 5.7.
Let $s$ be a $G$-invariant $L^{k}$-valued $(0,n_{-})$-forms on $M$. Then every $m\in M$ fulfilling $s(m)\neq 0$ lies in $M_{s}$.
Proof.
See the proofs of [9, Theorems 5.3 and 5.4], they hold true in our setting up to replacing the space of $G$-invariant holomorphic sections of $L^{k}$ with the space of $G$-invariant $L^{k}$-valued $(0,n_{-})$-forms. ∎
Before stating Theorem 5.9, we recall the following result from [11]. Although it can be deduced from the theory of Toeplitz operators discussed in the previous sections, here we present Theorem 5.8 in terms of the Bergman kernel, rather than the Szegő kernel. As we explained in Section 2, we identify the $G$-invariant Bergman projector $P^{(q),G}_{k}$ of $L^{k}$ with the $k$-th Fourier component of the $G$-invariant Szegő kernel $S^{(q)}_{G,k}$.
Let us premise a further piece of notation. We will identify the curvature form $R^{L}$ with the Hermitian matrix
$\dot{R}^{L}\in\mathcal{C}^{\infty}\bigl{(}M,\mathrm{End}\bigl{(}T^{1,0}M\bigr{)}\bigr{)},\qquad\big{\langle}R^{L}(z),U\wedge\overline{V}\big{\rangle}=\big{\langle}\dot{R}^{L}(z)U|V\big{\rangle},$
for every $U$ and $V$ in $T^{1,0}_{z}M$, $z\in M$. We denote by $W$ the subbundle of rank $q=n_{-}$ of $T^{1,0}M$ generated by the eigenvectors corresponding to negative eigenvalues of $\dot{R}^{L}$. Then, $\det\overline{W}^{*}:=\Lambda^{q}\overline{W}^{*}$ is a rank one subbundle, where $\overline{W}^{*}$ is the dual bundle of the complex conjugate bundle of $W$ and $\Lambda^{q}\overline{W}^{*}$ is the vector space of all finite sums $v_{1}\wedge\cdots\wedge v_{q}$, $v_{1},\dots,v_{q}\in\overline{W}^{*}$. We denote by $P_{\det\overline{W}^{*}}$ the orthogonal projection from $T^{*0,q}M$ onto $\det\overline{W}^{*}$.
Let $s$ and $s_{1}$ be local holomorphic trivializing sections of $L$ defined on open sets $D\subset M$ and ${D_{1}\subset M}$, respectively, $\left|s\right|^{2}_{h^{L}}={\rm e}^{-2\phi}$, $\left|s_{1}\right|^{2}_{h^{L}}={\rm e}^{-2\phi_{1}}$, $\phi\in\mathcal{C}^{\infty}(D,\mathbb{R})$, $\phi_{1}\in\mathcal{C}^{\infty}(D_{1},\mathbb{R})$ which can be assumed to be $G$-invariant. The localization of $P^{(q),G}_{k}$ with respect to $s$, $s_{1}$ is given by
$P^{(q),G}_{k,s,s_{1}}\colon\ \Omega^{0,q}_{c}(D)\rightarrow\Omega^{0,q}(D_{1}),\qquad u\rightarrow s^{-k}_{1}{\rm e}^{-k\phi_{1}}\bigl{(}P^{(q),G}_{k}\bigl{(}s^{k}{\rm e}^{k\phi}u\bigr{)}\bigr{)}.$
Let
$P^{(q),G}_{k,s,s_{1}}(x,y)\in\mathcal{C}^{\infty}\bigl{(}D_{1}\times D,T^{*0,q}M\boxtimes\bigl{(}T^{*0,q}M\bigr{)}^{*}\bigr{)}$
be the distribution kernel of $P^{(q),G}_{k,s,s_{1}}$. When $D=D_{1}$, $s=s_{1}$, we write $P^{(q),G}_{k,s}:=P^{(q),G}_{k,s,s}$, $P^{(q),G}_{k,s}(x,y):=P^{(q),G}_{k,s,s}(x,y)$.
Theorem 5.8.
With the notations and assumptions used above and recall that we let $q=n_{-}$. Let $s$ be a local holomorphic trivializing section of $L$ defined on an open set $D\subset M$, $\left|s\right|^{2}_{h^{L}}={\rm e}^{-2\phi}$.
Let $D$ be an open set of $X$ with $D\cap\Phi^{-1}(0)=\varnothing$. Then, as $k\rightarrow+\infty$, $P^{(q),G}_{k,s}=O(k^{-\infty})$ on $D$.
Let $m\in\Phi^{-1}(0)$ and let $U$ be an open set of $m$. Then, as $k\rightarrow+\infty$, in local coordinates defined in $U$ ((((as introduced in [11]),
$P^{(q),G}_{k,s}(x,y)={\rm e}^{{\rm i}k\Psi(x,y)}b(x,y,k)+O(k^{-\infty})\qquad\text{on}\ D\times D,$